# DevFD: Developmental Face Forgery Detection by Learning Shared and Orthogonal LoRA Subspaces

### v2

@@ 1 Introduction
The swift evolution of generative models, including Generative Adversarial Networks (GANs)  Oorloff and Yacoob (2023); Bounareli et al. (2023); Xu et al. (2022) and Diffusion models Kim et al. (2022); Xu et al. (2024); Kim et al. (2023), along with the rise of large models Zhu et al. (2023a); Chen et al. (2023); Li et al. (2023); Xue et al. (2024); Shi et al. (2025), propels the rapid advancement of face forgery technology, which introduces significant security risks and contributes to a crisis of public trust. These manipulation methods leave unique traces of forgery. Existing detectors perform well with known forgery types but struggle to identify novel ones. When new manipulation types arise, a trivial method is to add new forgery data to the primary dataset and retrain a binary model. However, the rapid evolution of forgery techniques necessitates a new learning strategy that can quickly adapt to new forgeries using limited computation and data.
Figure 1: (a) The t-SNE visualization of features extracted from the baseline on FF++ Rossler et al. (2019) and CDF2 Li et al. (2020b) shows that real faces exhibit a close distribution, while fake faces form five distinct clusters. This observation inspires us to (b) propose a developmental mixture of experts to model the continuous emergence of unknown fake faces using a set of orthogonal LoRA subspaces, while concurrently employing a dedicated LoRA to preserve the commonalities of authentic faces. A label-guided localized balancing strategy employs the LoRA sequence to separately model the common real faces and the incremental fake types information.
In contrast to the rapidly evolving fake faces, real faces possess distinct characteristics. Genuine facial data is relatively abundant, stable, and acquired through a single modality, such as camera imaging. In the context of forgery detection, it typically does not introduce bias. These properties, which remain invariant as forgery techniques evolve, endow real faces with commonalities that may easily be overlooked. Fig. 1(a) presents a simple t-SNE Van der Maaten and Hinton (2008) visualization experiment, using the baseline Chollet (2017) model to extract features from both the FF++ Rossler et al. (2019) and CDF2 Li et al. (2020b) datasets. The results reveal two key observations: (1) Real face samples from both datasets exhibit close distributions, while (2) forged samples form five distinct clusters. These observations inspire us to enable the model to adapt to emerging manipulation methods while learning the commonalities of real faces.
@@ 2.1 Face Forgery Detection
Face forgery detection requires the model to classify input images or videos as either real or fake. Related methods Chen et al. (2022); Dong et al. (2022); Liu et al. (2021); Wang and Ma (2024); Yan et al. (2025); Cheng et al. (2025) focus on identifying forgery cues present in the forged media and can be categorized into two major groups. The first category designs models that incorporate additional information to aid in detection, such as frequency analysis Qian et al. (2020); Frank et al. (2020); Li et al. (2021); Luo et al. (2021), wavelet Miao et al. (2023), graphics Zhu et al. (2021, 2023b), language descriptions Sun et al. (2023), and audio-video consistency Oorloff et al. (2024). The second category Li et al. (2020a); Shiohara and Yamasaki (2022); Larue et al. (2023) is data-driven, utilizing only real faces and training models through a self-supervised approach. With the continuous evolution of forgery methods, continual learning is applied to forgery detection. DFIL Pan et al. (2023) and HDP Sun et al. (2024) treat the learning of forgery datasets as a sequence of sub-tasks, requiring the model to learn sequentially across these tasks, achieving satisfactory performance. In this work, we design a Developmental MoE to address the challenges of continual learning in face forgery detection.
@@ 2.2 Continual Learning
Continual learning typically models the task as a sequence of sub-tasks, requiring the model to adapt to dynamic data distributions incrementally. The goal is to achieve high performance on newly learned tasks while avoiding catastrophic forgetting of information from previous tasks. Related approaches can be broadly divided into three categories: Regularization-based Approaches Ritter et al. (2018); Schwarz et al. (2018); Wang et al. (2021); Singh et al. (2021); Yan et al. (2021b); Zhang et al. (2021), Replay-based Approaches Simon et al. (2021); Riemer et al. (2018); Rebuffi et al. (2017b); Kurmi et al. (2021); Lopez-Paz and Ranzato (2017), and Optimization-based Approaches Lin et al. (2022); Kong et al. (2022); Liu and Liu (2022); Douillard et al. (2020); Pan et al. (2020). Among the latest methods, OrCo Ahmed et al. (2024) enhances generalization in class-incremental learning through orthogonality and contrast. Meanwhile, MoECL Yu et al. (2024) adapts to new tasks by introducing an MoE. DFIL Pan et al. (2023) and SUR-LID Cheng et al. (2024) achieve satisfactory results using replay mechanisms. In this work, we treat the iteration of fake faces as an incremental learning problem and employ a developmental architecture to facilitate continual learning.
@@ 2.3 Mixture of Experts
The Mixture of Experts (MoE) is a set of sparse models with multiple experts and a routing network. By training MoE in parallel, methods Mustafa et al. (2022); He et al. (2022); Zoph et al. (2022) enable efficient fine-tuning of large models with minimal computational cost. MoECL Yu et al. (2024) introduces MoE to continual learning, employing an intuitive trainable-freezing strategy to reduce the model’s forgetting rate. Low-rank Adaptation (LoRA) Hu et al. (2021) further decreases parameter and computational cost by mapping inputs into a low-rank subspace. O-LoRA Wang et al. (2023) utilizes LoRAs as experts and constrains the orthogonality of LoRA subspaces. InfLoRA Liang and Li (2024) further applies orthogonal LoRA to vision tasks, improving the performance of multiple sub-tasks through pre-defined orthogonal subspaces. MoEFFD Kong et al. (2024) integrates MoE and LoRA into face forgery detection, allowing each expert to specialize in extracting features related to different forgery techniques. We use the LoRA module to construct an MoE architecture to address continual face forgery detection.
@@ 3 Methodology
Figure 2: The proposed DevFD framework employs a Developmental MoE architecture to fine-tune the FFN layer in each transformer block. The architecture establishes a developmental LoRA sequence, which adds new branches as the number of tasks increases, enabling the model to handle emerging new types of forgeries. A new label-guided localized balancing strategy allocates the LoRA sequence into two purposes: the Real-LoRA refines the real face knowledge modeled by the backbone, while the Fake-LoRAs compose an orthogonal sequence to model the unique cues of fake faces. We integrate orthogonal gradients into the orthogonal loss to alleviate the interference of gradients on previously learned tasks during the training phase when orthogonality is not yet achieved, thereby achieving a lower rate of forgetting.
@@ 3.1 Developmental MoE Architecture
We propose a novel developmental MoE architecture using LoRAs as individual experts to fine-tune pre-trained models within different subspaces; the overview is shown in Fig. 2. First, we employ a Vision Transformer (ViT) pre-trained on real faces as the backbone, denoted as $g_{\mathbf{\Phi}}\left(f_{\mathbf{\Theta}}(\cdot)\right)$, where $f_{\mathbf{\Theta}}$ represents the pre-trained ViT network, and $g_{\mathbf{\Phi}}$ denotes the classifier. We freeze the parameter $\mathbf{\Theta}$ of the ViT backbone $f$ and introduce LoRA to fine-tune the Feed Forward Network (FFN) layer weights in every transformer block. For a parameter matrix of any FFN layer $\mathbf{W}\in\mathbb{R}^{d_{O}\times d_{I}}$ with input dimension $d_{I}$ and output dimension $d_{O}$, LoRA decomposes the parameter change matrix $\Delta W$ into the multiplication of two low-rank matrices:
@@
Here, $\mathbf{x}$ is the input, while $\mathbf{e}$ is the output. The LoRA modules in the sequence $\left\{\mathbf{A}_{j}\mathbf{B}_{j}\right\}_{i=0}^{t}$ are assigned by a label-guided localized balancing strategy (LLB) for two purposes: the Real-LoRA $\mathbf{A}_{0}\mathbf{B}_{0}$ is designed to learn and refine the common information of real faces modeled by the backbone, while the Fake-LoRAs $\left\{\mathbf{A}_{j}\mathbf{B}_{j}\right\}_{i=1}^{t}$ compose an orthogonal sequence for learning fake face information specific to each task.
@@ 3.2 Learning LoRA Subspaces enhanced by Orthogonal Gradients
The reason for catastrophic forgetting is that parameter adjustments while learning the current task disrupt the knowledge acquired from previous tasks Liang and Li (2024). The subspace established by LoRA can approximately represent the information acquired for a new task Wang et al. (2023). As a developmental structure, DevFD constrains the learning direction of the expanded branches for the new task to prevent interference with previously established subspaces by using an integrated orthogonal loss enhanced with orthogonal gradients. First, the learning of the t-th task can be modeled as follows:
@@
Why Orthogonal LoRA Subspaces are Still Forgetful? The adaptation of LoRA to new tasks involves fine-tuning the weight matrices within their respective subspaces. LoRA maps the input to the subspace : $\operatorname{span}\left\{\mathbf{b}_{1}^{t},\ldots,\mathbf{b}_{r}^{t}\right\}$, where the row vectors $\left\{\mathbf{b}_{1}^{t},\ldots,\mathbf{b}_{r}^{t}\right\}$ of the dimensionality reduction matrix $\mathbf{B}_{t}$ serve as a basis for this subspace. O-LoRA Wang et al. (2023) enforces the orthogonality between the basis $\left\{\mathbf{b}_{1}^{t},\ldots,\mathbf{b}_{r}^{t}\right\}$ and the basis $\left\{\mathbf{b}_{1}^{i},\ldots,\mathbf{b}_{r}^{i}\right\}$ of the forward LoRAs’ subspace, ensuring that the subspaces of the LoRA sequence are mutually orthogonal. Let $\mathbf{O_{i,t}}$ be the dot product of two bases of these subspaces. The orthogonality optimization objective can then be expressed as:
@@
Notably, the orthogonality condition exists in the form of an optimization objective. We will demonstrate that preventing the disruption of the information in the previous subspaces depends on the strict $\mathbf{B}_{t}^{T}\mathbf{B}_{i}=0$ throughout the training process. We begin with a proven proposition Liang and Li (2024): when the model learns the t-th task, fine-tuning $\mathbf{A}_{t}$ is equivalent to fine-tuning the original parameter matrix $\mathbf{W}$ within the subspace $\operatorname{span}\left\{\mathbf{b}_{1}^{t},\ldots,\mathbf{b}_{r}^{t}\right\}$, denoted as:
@@
Integrate Orthogonal Gradients into Orthogonal Loss. To address the issues discussed above, we choose to directly constrain the gradient space while optimizing the orthogonality of subspaces, resulting in a new integrated orthogonal loss. First, it is proven that the gradient update of a linear or convolution layer lies within the span of the input vectors Liang and Li (2023). We place the proofs in the Appendix B. This proposition allows us to transform the constraints on the gradient space into one on the input space. Let the input matrix be $\mathbf{H}\in\mathbb{R}^{d_{I}\times n}$, where $n$ is the batch size. Due to the inconsistency in matrix dimensions, we first perform Singular Value Decomposition (SVD) on $\mathbf{H}^{T}_{t}$:
@@ 3.3 Label-guided Localized Balancing Strategy
To model the commonalities of real faces and continuously refine the real face knowledge represented by the backbone in a continual learning sequence, we establish an additional LoRA Expert $\mathbf{A}_{0}\mathbf{B}_{0}$, referred to as Real-LoRA, and make it learnable across all tasks. However, the Real-LoRA introduces two issues. First, the additional Real-LoRA disrupts the orthogonality of the LoRA sequence, leading to interference with the Fake-LoRA sequence. Second, it cannot be guaranteed that the Real-LoRA will learn real-face information. To address these issues, we propose a label-guided localized balancing strategy. This strategy dynamically adjusts the responses of experts to different types of samples based on the labels during training. Specifically, it enables Real-LoRA to focus more on real faces, while Fake-LoRAs concentrate on learning forged faces in an orthogonal manner, preventing Real-LoRA from disrupting the orthogonality of Fake-LoRAs. The proposed strategy first takes the inputs from all LoRAs and sets a response matrix $\mathbf{I}\in\mathbb{R}^{t\times n}$, where an element $\mathbf{I}[k,l]$ indicates the response of the $k$-th expert to the $l$-th sample in the batch. It is defined as follows:
$\mathbf{I}_{k,l}=\sum_{j=1}^{T_{m}}\frac{\exp\left(\omega_{k}^{j}/\tau\right)}{\sum_{i=1}^{N}\exp\left(\omega_{i}^{j}/\tau\right)}.$ (10)
@@
Here, $\omega_{k}^{j}$ represents the output of the $k$-th expert for the $j$-th token in the $l$-th sample, $T_{m}$ denotes the number of tokens in the $l$-th sample, and $N$ indicates the total number of experts. All outputs of LoRA can be defined as a vector matrix $\mathbf{O}\in\mathbb{R}^{t\times n\times d_{o}}$, where each element is defined as the output of the $i$-th expert for the $j$-th sample, denoted as $\mathbf{O}[i,j]=\mathbf{A}_{i}\mathbf{B}_{i}\mathbf{H}[:,j]$. Based on the response matrix $\mathbf{I}$ and the output matrix $\mathbf{O}$, the forward process of the MoE can be represented as follows:
$\mathbf{e}=\mathbf{I^{T}}\mathbf{O}.$ (11)
@@
To allocate Real-LoRA and Fake-LoRAs for modeling the information of real and fake faces, respectively, we utilize labels to define a coefficient matrix $\mathbf{C}\in\mathbb{R}^{t\times n}$. An element $\mathbf{C}[i,j]$ represents the importance coefficient of the response at the corresponding position $\mathbf{I}[i,j]$:
@@
$\mathbf{C}_{i,j}=\left\{\begin{array}[]{ll}1+\delta,&\operatorname{Type}_{e}(i)=\operatorname{Type}_{h}(j)\\ 1-\delta,&\operatorname{Type}_{e}(i)\neq\operatorname{Type}_{h}(j)\end{array}\right$., (12)
@@
where $\operatorname{Type}_{e}(i)=\operatorname{Type}_{h}(j)$ indicates that the expert matches the category label corresponding to the sample, such as Real-LoRA for real faces, the coefficient matrix $\mathbf{C}$ increases the importance coefficient for this item; otherwise, $\mathbf{C}$ decreases the importance coefficient. The increment $\delta$ of the coefficient falls within the range of $[0,1]$. Based on the response matrix $\mathbf{I}$ and the coefficient matrix $\mathbf{C}$, the label-guided localized balancing loss constrains the dispersion of the weighted response matrix $\mathbf{I}\circ\mathbf{C}$:
$\mathcal{L}_{\text{llb}}=\frac{\sigma^{2}(\mathbf{\mathbf{I}\circ\mathbf{C}})}{\mu(\mathbf{\mathbf{I}\circ\mathbf{C}})},$ (13)
@@
where $\sigma^{2}(\cdot)$ and $\mu(\cdot)$ represent the variance and mean, respectively, and $\circ$ is element-wise multiplication. By optimizing this loss function, the LoRA that matches the sample type will yield a greater response for that class of samples, thereby learning more knowledge from them. Additionally, this router does not select specific LoRAs but encourages collaboration among all LoRAs.
@@
When the model begins to learn a new task, DevFD establishes a new LoRA branch and freeze the other Fake-LoRAs in the orthogonal sequence $\left\{\mathbf{A}_{j}\mathbf{B}_{j}\right\}_{i=1}^{t}$, allowing only the newly added LoRA $\mathbf{A}_{t}\mathbf{B}_{t}$ and the Real-LoRA $\mathbf{A}_{0}\mathbf{B}_{0}$ to be trainable. During inference, it aggregates the outputs of all LoRA experts to make a joint decision, fostering collaboration among them.

### v3

@@ 1 Introduction
The swift evolution of generative models, including Generative Adversarial Networks (GANs)  gan_gen1; gan_gen2; gan_gen3 and Diffusion models diffface; diff2; diff3, along with the rise of large models minigpt; minigptv2; blip2; blip3; shieldllm, propels the rapid advancement of face forgery technology, which introduces significant security risks and contributes to a crisis of public trust. These manipulation methods leave unique traces of forgery. Existing detectors perform well with known forgery types but struggle to identify novel ones. When new manipulation types arise, a trivial method is to add new forgery data to the primary dataset and retrain a binary model. However, the rapid evolution of forgery techniques necessitates a new learning strategy that can quickly adapt to new forgeries using limited computation and data.
Figure 1: (a) The t-SNE visualization of features extracted from the baseline on FF++ ffpp and CDF2 celeb shows that real faces exhibit a close distribution, while fake faces form five distinct clusters. This observation inspires us to (b) propose a developmental mixture of experts to model the continuous emergence of unknown fake faces using a set of orthogonal LoRA subspaces, while concurrently employing a dedicated LoRA to preserve the commonalities of authentic faces. A label-guided localized balancing strategy employs the LoRA sequence to separately model the common real faces and the incremental fake types information.
In contrast to the rapidly evolving fake faces, real faces possess distinct characteristics. Genuine facial data is relatively abundant, stable, and acquired through a single modality, such as camera imaging. In the context of forgery detection, it typically does not introduce bias. These properties, which remain invariant as forgery techniques evolve, endow real faces with commonalities that may easily be overlooked. Fig. 1(a) presents a simple t-SNE tsne visualization experiment, using the baseline xception model to extract features from both the FF++ ffpp and CDF2 celeb datasets. The results reveal two key observations: (1) Real face samples from both datasets exhibit close distributions, while (2) forged samples form five distinct clusters. These observations inspire us to enable the model to adapt to emerging manipulation methods while learning the commonalities of real faces.
@@ 2.1 Face Forgery Detection
Face forgery detection requires the model to classify input images or videos as either real or fake. Related methods det1; det2; det3; detllm; vbsta; clformer focus on identifying forgery cues present in the forged media and can be categorized into two major groups. The first category designs models that incorporate additional information to aid in detection, such as frequency analysis f3net; freq1; freq2; freq3, wavelet wavelet, graphics fd2net; fd2net2, language descriptions langu, and audio-video consistency avff. The second category fxray; sbi; seeable is data-driven, utilizing only real faces and training models through a self-supervised approach. With the continuous evolution of forgery methods, continual learning is applied to forgery detection. DFIL dfil and HDP hdp treat the learning of forgery datasets as a sequence of sub-tasks, requiring the model to learn sequentially across these tasks, achieving satisfactory performance. In this work, we design a Developmental MoE to address the challenges of continual learning in face forgery detection.
@@ 2.2 Continual Learning
Continual learning typically models the task as a sequence of sub-tasks, requiring the model to adapt to dynamic data distributions incrementally. The goal is to achieve high performance on newly learned tasks while avoiding catastrophic forgetting of information from previous tasks. Related approaches can be broadly divided into three categories: Regularization-based Approaches reg1; reg2; reg3; reg4; reg5; reg6, Replay-based Approaches replay1; replay2; replay3; replay4; replay5, and Optimization-based Approaches opt1; opt2; opt3; opt4; opt5. Among the latest methods, OrCo orco enhances generalization in class-incremental learning through orthogonality and contrast. Meanwhile, MoECL moecl adapts to new tasks by introducing an MoE. DFIL dfil and SUR-LID surlid achieve satisfactory results using replay mechanisms. In this work, we treat the iteration of fake faces as an incremental learning problem and employ a developmental architecture to facilitate continual learning.
@@ 2.3 Mixture of Experts
The Mixture of Experts (MoE) is a set of sparse models with multiple experts and a routing network. By training MoE in parallel, methods limoe; fastermoe; stmoe enable efficient fine-tuning of large models with minimal computational cost. MoECL moecl introduces MoE to continual learning, employing an intuitive trainable-freezing strategy to reduce the model’s forgetting rate. Low-rank Adaptation (LoRA) lora further decreases parameter and computational cost by mapping inputs into a low-rank subspace. O-LoRA olora utilizes LoRAs as experts and constrains the orthogonality of LoRA subspaces. InfLoRA inflora further applies orthogonal LoRA to vision tasks, improving the performance of multiple sub-tasks through pre-defined orthogonal subspaces. MoEFFD moeffd integrates MoE and LoRA into face forgery detection, allowing each expert to specialize in extracting features related to different forgery techniques. We use the LoRA module to construct an MoE architecture to address continual face forgery detection.
@@ 3 Methodology
Figure 2: The proposed DevFD framework employs a Developmental MoE architecture to fine-tune the FFN layer in each transformer block. The architecture establishes a developmental LoRA sequence, which adds new branches as the number of tasks increases, enabling the model to handle emerging new types of forgeries. A new label-guided localized balancing strategy allocates the LoRA sequence into two purposes: the Real-LoRA fine-tune and refine knowledge about real faces, while the Fake-LoRAs compose an orthogonal sequence to model the unique cues of fake faces. We integrate orthogonal gradients into the orthogonal loss to alleviate the interference of gradients on previously learned tasks during the training phase when orthogonality is not yet achieved, thereby achieving a lower rate of forgetting.
@@ 3.1 Developmental MoE Architecture
We propose a novel developmental MoE architecture using LoRAs as individual experts to fine-tune pre-trained models within different subspaces; the overview is shown in Fig. 2. First, we employ a Vision Transformer (ViT) pre-trained on real faces as the backbone, denoted as $g_{\mathbf{\Phi}}\left(f_{\mathbf{\Theta}}(\cdot)\right)$, where $f_{\mathbf{\Theta}}$ represents the pre-trained ViT network, and $g_{\mathbf{\Phi}}$ denotes the classifier. We freeze the parameter $\mathbf{\Theta}$ of the ViT backbone $f$ and introduce LoRA to fine-tune the Feed Forward Network (FFN) weights in each transformer block. For a parameter matrix of any FFN layer $\mathbf{W}\in\mathbb{R}^{d_{O}\times d_{I}}$ with input dimension $d_{I}$ and output dimension $d_{O}$, LoRA decomposes the parameter change matrix $\Delta W$ into the multiplication of two low-rank matrices:
@@
Here, $\mathbf{x}$ is the input, while $\mathbf{e}$ is the output. The LoRA modules in the sequence $\left\{\mathbf{A}_{j}\mathbf{B}_{j}\right\}_{j=0}^{t}$ are assigned by a label-guided localized balancing strategy (LLB) for two purposes: the Real-LoRA $\mathbf{A}_{0}\mathbf{B}_{0}$ is designed to learn and refine the common information of real faces modeled by the backbone, while the Fake-LoRAs $\left\{\mathbf{A}_{j}\mathbf{B}_{j}\right\}_{j=1}^{t}$ compose an orthogonal sequence for learning fake face information specific to each task.
@@ 3.2 Learning LoRA Subspaces enhanced by Orthogonal Gradients
The reason for catastrophic forgetting is that parameter adjustments while learning the current task disrupt the knowledge acquired from previous tasks inflora. The subspace established by LoRA can approximately represent the information acquired for a new task olora. As a developmental structure, DevFD constrains the learning direction of the expanded branches for the new task to prevent interference with previously established subspaces by using an integrated orthogonal loss enhanced with orthogonal gradients. First, the learning of the t-th task can be modeled as follows:
@@
Why Orthogonal LoRA Subspaces are Still Forgetful? The adaptation of LoRA to new tasks involves fine-tuning the weight matrices within their respective subspaces. LoRA maps the input to the subspace : $\operatorname{span}\left\{\mathbf{b}_{1}^{t},\ldots,\mathbf{b}_{r}^{t}\right\}$, where the row vectors $\left\{\mathbf{b}_{1}^{t},\ldots,\mathbf{b}_{r}^{t}\right\}$ of the dimensionality reduction matrix $\mathbf{B}_{t}$ serve as a basis for this subspace. O-LoRA olora enforces the orthogonality between the basis $\left\{\mathbf{b}_{1}^{t},\ldots,\mathbf{b}_{r}^{t}\right\}$ and the basis $\left\{\mathbf{b}_{1}^{i},\ldots,\mathbf{b}_{r}^{i}\right\}$ of the forward LoRAs’ subspace, ensuring that the subspaces of the LoRA sequence are mutually orthogonal. Let $\mathbf{O_{i,t}}$ be the dot product of two bases of these subspaces. The orthogonality optimization objective can then be expressed as:
@@
Notably, the orthogonality condition exists in the form of an optimization objective. We will demonstrate that preventing the disruption of the information in the previous subspaces depends on the strict $\mathbf{B}_{t}^{T}\mathbf{B}_{i}=0$ throughout the training process. We begin with a proven proposition inflora: when the model learns the t-th task, fine-tuning $\mathbf{A}_{t}$ is equivalent to fine-tuning the original parameter matrix $\mathbf{W}$ within the subspace $\operatorname{span}\left\{\mathbf{b}_{1}^{t},\ldots,\mathbf{b}_{r}^{t}\right\}$, denoted as:
@@
Integrate Orthogonal Gradients into Orthogonal Loss. To address the issues discussed above, we choose to directly constrain the gradient space while optimizing the orthogonality of subspaces, resulting in a new integrated orthogonal loss. First, it is proven that the gradient update of a linear or convolution layer lies within the span of the input vectors api. We place the proofs in the Appendix B. This proposition allows us to transform the constraints on the gradient space into one on the input space. Let the input matrix be $\mathbf{H}\in\mathbb{R}^{d_{I}\times n}$, where $n$ is the batch size. Due to the inconsistency in matrix dimensions, we first perform Singular Value Decomposition (SVD) on $\mathbf{H}^{T}_{t}$:
@@ 3.3 Label-guided Localized Balancing Strategy
To model the commonalities of real faces and continuously refine the real face knowledge represented by the backbone in a continual learning sequence, we establish an additional LoRA Expert $\mathbf{A}_{0}\mathbf{B}_{0}$, referred to as Real-LoRA, and make it learnable across all tasks. However, the Real-LoRA introduces two issues. First, the additional Real-LoRA disrupts the orthogonality of the LoRA sequence, leading to interference with the Fake-LoRA sequence. Second, it cannot be guaranteed that the Real-LoRA will learn real-face information. To address these issues, we propose a label-guided localized balancing strategy. This strategy dynamically adjusts the responses of experts to different types of samples based on the labels during training. Specifically, it enables Real-LoRA to focus more on real faces, while Fake-LoRAs concentrate on learning forged faces in an orthogonal manner, preventing Real-LoRA from disrupting the orthogonality of Fake-LoRAs. The proposed strategy first takes the inputs from all LoRAs and sets a response matrix $\mathbf{I}\in\mathbb{R}^{(t+1)\times n}$, where $t+1$ indicates the total number of experts. An element $\mathbf{I}_{k,l}$ indicates the response of the $k$-th expert to the $l$-th sample in the batch. It is defined as follows:
$\mathbf{I}_{k,l}=\sum_{j=1}^{T_{m}}\frac{\exp\left(\omega_{k,l}^{j}/\tau\right)}{\sum_{i=0}^{t}\exp\left(\omega_{i,l}^{j}/\tau\right)}.$ (10)
@@
Here, $\omega_{k,l}^{j}$ represents the output of the $k$-th expert for the $j$-th token in the $l$-th sample, $T_{m}$ denotes the number of tokens in the $l$-th sample. All outputs of LoRA can be defined as a vector matrix $\mathbf{O}\in\mathbb{R}^{(t+1)\times n\times d_{o}}$, where each element is defined as the output of the $k$-th expert for the $l$-th sample. Based on the response matrix $\mathbf{I}$ and the output matrix $\mathbf{O}$, the forward process computes the final output $\mathbf{e}\in\mathbb{R}^{n\times d_{o}}$, where the output vector $\mathbf{e}_{l}\in\mathbb{R}^{d_{o}}$ for the $l$-th sample is defined as:
$\mathbf{e}_{l}=\sum_{k=0}^{t}\mathbf{I}_{k,l}\cdot\mathbf{O}_{k,l}$ (11)
@@
To allocate Real-LoRA and Fake-LoRAs for modeling the information of real and fake faces, respectively, we utilize labels to define a balancing coefficient matrix $\mathbf{C}\in\mathbb{R}^{(t+1)\times n}$. An element $\mathbf{C}_{k,l}$ represents the balancing weight applied to the response at the corresponding position $\mathbf{I}_{k,l}$:
@@
$\mathbf{C}_{k,l}=\left\{\begin{array}[]{ll}1-\delta,&\operatorname{Type}_{e}(k)=\operatorname{Type}_{h}(l)\\ 1+\delta,&\operatorname{Type}_{e}(k)\neq\operatorname{Type}_{h}(l)\end{array}\right$., (12)
@@
where $\operatorname{Type}_{e}(k)=\operatorname{Type}_{h}(l)$ indicates that the expert matches the category label corresponding to the sample, such as Real-LoRA for real faces, the coefficient matrix $\mathbf{C}$ decreases the balancing coefficient for this item; otherwise, $\mathbf{C}$ increases the balancing coefficient. The increment $\delta$ of the coefficient falls within the range of $[0,1]$. Based on the response matrix $\mathbf{I}$ and the balancing coefficient matrix $\mathbf{C}$, the LLB loss constrains the dispersion of the weighted response matrix $\mathbf{I}\circ\mathbf{C}$:
$\mathcal{L}_{\text{llb}}=\frac{\sigma^{2}(\mathbf{I}\circ\mathbf{C})}{\mu(\mathbf{I}\circ\mathbf{C})},$ (13)
@@
where $\sigma^{2}(\cdot)$ and $\mu(\cdot)$ represent the variance and mean, respectively, and $\circ$ is element-wise multiplication. This loss function forces the weighted responses ($\mathbf{I}\circ\mathbf{C}$) to become as balanced as possible. By optimizing this loss function, the model is thus compelled to generate a greater original response $\mathbf{I}_{k,l}$ for the matching pairs (to compensate for their small $1-\delta$ balancing coefficients) and a smaller original response for the mismatched pairs. Additionally, this router does not select specific LoRAs but encourages collaboration among all LoRAs.
@@
When the model begins to learn a new task, DevFD establishes a new LoRA branch and freeze the other Fake-LoRAs in the orthogonal sequence $\left\{\mathbf{A}_{j}\mathbf{B}_{j}\right\}_{j=1}^{t}$, allowing only the newly added LoRA $\mathbf{A}_{t}\mathbf{B}_{t}$ and the Real-LoRA $\mathbf{A}_{0}\mathbf{B}_{0}$ to be trainable. During inference, it aggregates the outputs of all LoRA experts to make a joint decision, fostering collaboration among them.